# Early Galaxies and Early Dark Energy: A Unified Solution to the Hubble Tension and Puzzles of Massive Bright Galaxies revealed by JWST

### v1

@@ 4.1 Benchmark the model at $4\lesssim z\lesssim 9$
As the first step, we calibrate our model based on the observational constraints of UVLFs at $4\lesssim z\lesssim 9$. In this paper, we include the HST observations compiled in Vogelsberger et al. (2020) and those from McLeod et al. (2016); Oesch et al. (2018); Morishita et al. (2018); Stefanon et al. (2019); Bowler et al. (2020); Bouwens et al. (2021). For JWST constraints, we include constraints based on photometrically selected galaxies from Castellano et al. (2022); Finkelstein et al. (2022); Naidu et al. (2022b); Adams et al. (2023a); Bouwens et al. (2023b); Bouwens et al. (2023a); Donnan et al. (2023); Harikane et al. (2023); Leethochawalit et al. (2023); Morishita & Stiavelli (2023); Pérez-González et al. (2023); Robertson et al. (2023b); McLeod et al. (2024); Donnan et al. (2024); Casey et al. (2024). Furthermore, we include constraints based only on spectroscopically-confirmed galaxies (Harikane et al., 2024, see the references therein).
@@ 4.3 Explore the model parameter space
To illustrate various model variations to reconcile JWST results with theoretical models, Figure 7 examines the parameter space of UV variability and the median UV radiation yield in the $\Lambda$CDM and EDE cosmology. For results based on JWST spectroscopy, we adopt the lower limit estimated in Harikane et al. (2024). For photometric constraints, we consider the model to be acceptable when $\log_{10}{\Phi}(M_{\rm UV}=-20)>-4.8$ at $z\simeq 12$, and $\log_{10}{\Phi}(M_{\rm UV}=-19)>-4.4$ and $-4.5$ at $z=$14 and 16. These are fairly rough estimates (i.e. without a quantitative measure of fitting residuals), but we have explicitly examined the UVLF predictions around these threshold values and find that they capture the goodness of fit reasonably well. We scan the parameter space by modifying UV variability ($A(z)$ or equivalently $\sigma_{\rm UV}(10^{10.5}{\,\rm M_{\odot}})$) and the normalization of the SFE in our model, and identify the regime where theoretically predicted UV bright galaxy abundance exceeds the observed values. We also show the impact of IMF variations as in Shen et al. (2023), where a Salpeter IMF leads to about $60\%$ increase in $\kappa_{\rm UV}$ (Madau & Dickinson, 2014) and an extremely top-heavy IMF (e.g. Inayoshi et al., 2022) leads to $\kappa_{\rm UV}$ dropping by $55\%$.
@@
Qualitative differences show up when confronting the $z\simeq 16$ constraints. In $\Lambda$CDM, an extremely high, constant $\sigma_{\rm UV}\simeq 2.2$ mag is required, which would imply a strong non-linear transition of UV variability at $z\gtrsim 14$. In the EDE cosmology, it becomes feasible with a suggested $\sigma_{\rm UV}(10^{10.5}{\,\rm M_{\odot}})\simeq 1.3$ mag. One could also achieve this by boosting the SFE by a factor of $\sim 5$ and the implied $\epsilon_{\ast}$ at $\sim 10^{10.5}{\,\rm M_{\odot}}$ will be $\sim 10\%$, which could be reached in the FFB scenario (Li et al., 2023). The introduction of EDE eases the level of tuning to the galaxy formation model to reconcile these extreme JWST results. However, we are cautious in making strong conclusions about $\Lambda$CDM based on the comparison at this redshift, since the constraints here are largely based on a handful of photometrically-selected galaxies (e.g. Bouwens et al., 2023b; Harikane et al., 2023) and remain uncertain.
Figure 8: Galaxy UVLF at $z\simeq 14$ in the following four scenarios: (1) the constant UV variability model considered in Shen et al. (2023); (2) boosting the SFE by a factor of three while assuming the benchmark $\sigma_{\rm UV}(M_{\rm halo})$ in $\Lambda$CDM; (3) adopt appropriate modifications to $\sigma_{\rm UV}(M_{\rm halo})$ in $\Lambda$CDM and (4) EDE, respectively. The observed UVLF at $z\simeq 14$ are equally well explained by these adjustments to the galaxy formation model.
@@ 4.4 Host haloes of UV-bright galaxies at cosmic dawn
In Figure 8, we show the UVLFs at $z\simeq 14$ in four different scenarios that could equally well match observational constraints. This includes the model with constant $\sigma_{\rm UV}$ as explored in Shen et al. (2023), the approach of boosting the SFE by three times with respect to the benchmark model at low redshifts (see similar values found in e.g. Yung et al. 2024b), and the approach of adjusting the normalizations of $\sigma_{\rm UV}(M_{\rm halo})$ in the $\Lambda$CDM and EDE cosmologies. However, due to different levels of $\sigma_{\rm UV}$ adopted, the host halo mass of observed UV-bright galaxies can be rather different in these scenarios. In the top panel of Figure 9, we show the median host halo mass of galaxies at a given $M_{\rm UV}$ at $z\simeq 14$. Assuming only varying UV variability, the typical halo mass of observed galaxies in $\Lambda$CDM is around $10^{9}{\,\rm M_{\odot}}$ at $z\simeq 14$, the mass scale of current-day ultra-faint dwarfs. The descendants of these haloes will end up with $\sim 10^{13}$ $10^{14}{\,\rm M_{\odot}}$ when evolving to $z=0$ (e.g. Lu et al., 2024). However, in the EDE cosmology, due to the much smaller $\sigma_{\rm UV}$ required, the host haloes are about $0.5$ $1$ dex heavier than in $\Lambda$CDM. It is even more of the case if one attempts to solve the tension by only adjusting SFE. We find that the host halo mass in this scenario will be around $10^{10.5}{\,\rm M_{\odot}}$.
@@
Figure 10: Cosmic stellar mass density for galaxies with stellar mass above $M_{\ast}$ at $z\simeq 8$, 9, and 12. The dashed lines show the upper limit derived by converting all the baryons in the Universe to stars in the assumed cosmology. The solid lines show predictions from our benchmark model at $z\simeq 8,$ 9, and the slightly tuned model at $z\simeq 12$ which match the UVLF constraints in Figure 6. The dotted and dotted-dashed lines show two alternative scenarios in EDE when (1) $\sigma_{\rm UV}(10^{10.5}{\,\rm M_{\odot}})$ is further increased to $1.3$ mag (the value needed for the $z\simeq 16$ UVLF constraints) or (2) SFE is enhanced by a factor of five. We compare them to the observational constraints from JWST reported in Labbé et al. (2023); Akins et al. (2023); Casey et al. (2024), and Wang et al. (2024a) (“medium” estimate as the circle, “maximum” as the star). Desprez et al. (2024) revisited the CEERS sample used in Labbé et al. (2023) and derived alternative constraints when the single most massive outlier was removed. We also show estimates by integrating the Schechter function fits of stellar mass functions from Stefanon et al. (2021) based on HST+IRAC and Weibel et al. (2024) based on JWST. IMF-related corrections on stellar masses are made for all these data. The most extreme observational constraints from JWST, which are in conflict with $\Lambda$CDM, can be safely accommodated in the EDE cosmology. The benchmark model predictions at $z\simeq 8$ and 9 marginally agree with some less stringent observational constraints. At $z\simeq 12$, a boosted SFE can better explain the large cosmic stellar mass densities compared to enhanced UV variabilities.
@@ 4.5 Implications for the overly-massive galaxies
where $M_{\rm min}=10^{8}{\,\rm M_{\odot}}$ is the minimum halo mass we start from (the SFE below which is assumed to be a constant) which does not affect our result, $R$ is the mass return fraction of stars taken to be $0.1$ for the young and low-metallicity stellar populations at high redshifts (e.g. Hopkins et al., 2023, Feldmann in prep.), $\mathcal{F}(\sigma_{\rm sf})$ is the factor considering the difference between mean SFR and median SFR with the logarithm scatter $\sigma_{\rm sfr}$ in unit of dex
@@
In Figure 10, we compare the stellar mass density in $\Lambda$CDM and EDE with observational constraints from JWST. We show the model predictions at $z\simeq 8$, 9, and 12, which corresponds to the constraints from CEERS (Labbé et al., 2023), COSMOS-Web (Casey et al., 2024), RUBIES (Wang et al., 2024a), and Akins et al. (2023). For reference, we also show estimates by integrating the Schechter function fits of stellar mass functions from Stefanon et al. (2021) and Weibel et al. (2024). A notable tension emerges with the Labbé et al. (2023) results surpassing the maximum achievable stellar mass density in the $\Lambda$CDM framework. However, EDE exhibits no such challenge, with all data points comfortably below the maximum threshold at the 1-$\sigma$ level. It is feasible to produce these massive galaxies with a physically realistic SFE. The stellar mass density produced by our fiducial star-formation model (with $A(z)$ chosen to match the UVLF at the corresponding redshift) in EDE is marginally consistent with findings from Wang et al. (2024a) and Desprez et al. (2024) at $z\simeq 8$ (who revisited the CEERS sample and tested removing the single most massive outlier), and the Cosmos-Web results at $z\sim 10$ (Casey et al., 2024). However, discrepancies persist at $z\simeq 12$. We explore two potential model variations, increasing $A(z)$ such that $\sigma_{\rm UV}(10^{10.5})\simeq 1.3$ mag or increasing the SFE by five times. As shown in Figure 7, these two can make the UVLF at $z\simeq 16$ consistent with observations. Since the impact of UV variability on $\rho_{\ast}$ is primarily through $\mathcal{F}(\sigma_{\rm sfr})$, it is very hard to reconcile observations purely on the UV variability direction. Alternatively, a heightened SFE emerges as a potential solution for reconciling the model with observational discrepancies in this context, as it not only increases the $M_{\ast}$ at a given $M_{\rm halo}$ but also lowers the integration range in Equation 10.
@@ Acknowledgements
We thank Robert Feldmann for suggestions that improved the manuscript. MV acknowledges support through NASA ATP 19-ATP19-0019, 19-ATP19-0020, 19-ATP19-0167, and NSF grants AST-1814053, AST-1814259, AST-1909831, AST-2007355 and AST-2107724. MBK acknowledges support from NSF CAREER award AST-1752913, NSF grants AST-1910346 and AST-2108962, NASA grant 80NSSC22K0827, and HST-GO-16686, HST-AR-17028, and HST-AR-17043 from the Space Telescope Science Institute (STScI), which is operated by AURA, Inc., under NASA contract NAS5-26555. RPN acknowledges support for this work provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by STScI.

### v2

@@ 4.1 Benchmark the model at $4\lesssim z\lesssim 9$
As the first step, we calibrate our model based on the observational constraints of UVLFs at $4\lesssim z\lesssim 9$. In this paper, we include the HST observations compiled in Vogelsberger et al. (2020) and those from McLeod et al. (2016); Oesch et al. (2018); Morishita et al. (2018); Stefanon et al. (2019); Bowler et al. (2020); Bouwens et al. (2021). For JWST constraints, we include constraints based on photometrically selected galaxies from Castellano et al. (2022); Finkelstein et al. (2022); Naidu et al. (2022b); Adams et al. (2023a); Bouwens et al. (2023b); Bouwens et al. (2023a); Donnan et al. (2023); Harikane et al. (2023); Leethochawalit et al. (2023); Morishita & Stiavelli (2023); Pérez-González et al. (2023); Robertson et al. (2023b); McLeod et al. (2024); Donnan et al. (2024); Casey et al. (2024). Furthermore, we include constraints based only on spectroscopically-confirmed galaxies (Harikane et al., 2024b, a, see the references therein).
@@ 4.3 Explore the model parameter space
To illustrate various model variations to reconcile JWST results with theoretical models, Figure 7 examines the parameter space of UV variability and the median UV radiation yield in the $\Lambda$CDM and EDE cosmology. For results based on JWST spectroscopy, we adopt the lower limit estimated in Harikane et al. (2024b, a) at $z\simeq 12$. For photometric constraints, we consider the model to be acceptable when $\log_{10}{\Phi}(M_{\rm UV}=-20)>-4.8$ at $z\simeq 12$, and $\log_{10}{\Phi}(M_{\rm UV}=-19)>-4.4$ and $-4.5$ at $z=$14 and 16. These are fairly rough estimates (i.e. without a quantitative measure of fitting residuals), but we have explicitly examined the UVLF predictions around these threshold values and find that they capture the goodness of fit reasonably well. We scan the parameter space by modifying UV variability ($A(z)$ or equivalently $\sigma_{\rm UV}(10^{10.5}{\,\rm M_{\odot}})$) and the normalization of the SFE in our model, and identify the regime where theoretically predicted UV bright galaxy abundance exceeds the observed values. We also show the impact of IMF variations as in Shen et al. (2023), where a Salpeter IMF leads to about $60\%$ increase in $\kappa_{\rm UV}$ (Madau & Dickinson, 2014) and an extremely top-heavy IMF (e.g. Inayoshi et al., 2022) leads to $\kappa_{\rm UV}$ dropping by $55\%$.
@@
Qualitative differences show up when confronting the $z\simeq 16$ constraints. In $\Lambda$CDM, an extremely high, constant $\sigma_{\rm UV}\simeq 2.2$ mag is required, which would imply a strong non-linear transition of UV variability at $z\gtrsim 14$. In the EDE cosmology, it becomes feasible with a suggested $\sigma_{\rm UV}(10^{10.5}{\,\rm M_{\odot}})\simeq 1.3$ mag. One could also achieve this by boosting the SFE by a factor of $\sim 5$ and the implied $\epsilon_{\ast}$ at $\sim 10^{10.5}{\,\rm M_{\odot}}$ will be $\sim 10\%$, which is below the maximum value found in literature and could be easily reached in the FFB scenario (Li et al., 2023). The introduction of EDE eases the level of tuning to the galaxy formation model to reconcile these extreme JWST results. However, we are cautious in making strong conclusions about $\Lambda$CDM based on the comparison at this redshift, since the constraints here are largely based on a handful of photometrically-selected galaxies (e.g. Bouwens et al., 2023b; Harikane et al., 2023) and remain uncertain.
Figure 8: Galaxy UVLF at $z\simeq 14$ in the following four scenarios: (1) the constant UV variability model considered in Shen et al. (2023); (2) boosting the SFE by a factor of three while assuming the benchmark $\sigma_{\rm UV}(M_{\rm halo})$ in $\Lambda$CDM; (3) adopt appropriate modifications to the normalization of $\sigma_{\rm UV}(M_{\rm halo})$ in $\Lambda$CDM and (4) EDE, respectively. The observed UVLF at $z\simeq 14$ are equally well explained by these adjustments to the galaxy formation model.
@@ 4.4 Host haloes of UV-bright galaxies at cosmic dawn
In Figure 8, we show the UVLFs at $z\simeq 14$ in four different scenarios that could equally well match observational constraints. This includes the model with constant $\sigma_{\rm UV}$ as explored in Shen et al. (2023), the approach of boosting the SFE by three times with respect to the benchmark model at low redshifts (see similar values found in e.g. Yung et al. 2024b), and the approach of adjusting the normalizations of $\sigma_{\rm UV}(M_{\rm halo})$ in the $\Lambda$CDM and EDE cosmologies as discussed above. However, due to different levels of $\sigma_{\rm UV}$ adopted, the host halo mass of observed UV-bright galaxies can be rather different in these scenarios. In the top panel of Figure 9, we show the median host halo mass of galaxies at a given $M_{\rm UV}$ at $z\simeq 14$. Assuming only varying UV variability, the typical halo mass of observed galaxies in $\Lambda$CDM is around $10^{9}{\,\rm M_{\odot}}$ at $z\simeq 14$, the mass scale of current-day ultra-faint dwarfs. The descendants of these haloes will end up with $\sim 10^{13}$ $10^{14}{\,\rm M_{\odot}}$ when evolving to $z=0$ (e.g. Lu et al., 2024). However, in the EDE cosmology, due to the much smaller $\sigma_{\rm UV}$ required, the host haloes are about $0.5$ $1$ dex heavier than in $\Lambda$CDM. It is even more of the case if one attempts to solve the tension by only adjusting SFE. We find that the host halo mass in this scenario will be around $10^{10.5}{\,\rm M_{\odot}}$.
@@
Figure 10: Cosmic stellar mass density for galaxies with stellar mass above $M_{\ast}$ at $z\simeq 8$, 9, and 12. The dashed lines show the upper limit derived by converting all the baryons in the Universe to stars in the assumed cosmology. The solid lines show predictions from our benchmark model at $z\simeq 8,$ 9, and the slightly tuned model at $z\simeq 12$ which match the UVLF constraints in Figure 6. The dotted and dotted-dashed lines show two alternative scenarios in EDE when (1) $\sigma_{\rm UV}(10^{10.5}{\,\rm M_{\odot}})$ is further increased to $1.3$ mag (the value needed for the $z\simeq 16$ UVLF constraints) or (2) SFE is enhanced by a factor of five. We compare them to the observational constraints from JWST reported in Labbé et al. (2023); Akins et al. (2023); Casey et al. (2024), and Wang et al. (2024a) (“medium” estimate as the circle, “maximum” as the star). Desprez et al. (2024) revisited the CEERS sample used in Labbé et al. (2023) and derived alternative constraints when the single most massive outlier was removed. We also show estimates by integrating the Schechter function fits of stellar mass functions from Stefanon et al. (2021); Weibel et al. (2024); Harvey et al. (2024). IMF-related corrections on stellar masses are made for all these data. The most extreme observational constraints from JWST, which are in conflict with $\Lambda$CDM, can be safely accommodated in the EDE cosmology. The benchmark model predictions at $z\simeq 8$ and 9 marginally agree with some less stringent observational constraints. At $z\simeq 12$, a boosted SFE can better explain the large cosmic stellar mass densities compared to enhanced UV variabilities.
@@ 4.5 Implications for the overly-massive galaxies
where $M_{\rm min}=10^{8}{\,\rm M_{\odot}}$ is the minimum halo mass we start from (the SFE below which is assumed to be a constant) which does not affect our result, $R$ is the mass return fraction of stars taken to be $0.1$ for the young and low-metallicity stellar populations at high redshifts (e.g. Hopkins et al., 2018, 2023; Feldmann et al., 2024), $\mathcal{F}(\sigma_{\rm sf})$ is the factor considering the difference between mean SFR and median SFR with the logarithm scatter $\sigma_{\rm sfr}$ in unit of dex
@@
In Figure 10, we compare the stellar mass density in $\Lambda$CDM and EDE with observational constraints from JWST. We show the model predictions at $z\simeq 8$, 9, and 12, which corresponds to the constraints from CEERS (Labbé et al., 2023), COSMOS-Web (Casey et al., 2024), RUBIES (Wang et al., 2024a), and Akins et al. (2023). For reference, we also show estimates by integrating the Schechter function fits of stellar mass functions from Stefanon et al. (2021); Weibel et al. (2024) and Harvey et al. (2024). A notable tension emerges with the Labbé et al. (2023) results surpassing the maximum achievable stellar mass density in the $\Lambda$CDM framework. However, EDE exhibits no such challenge, with all data points comfortably below the maximum threshold at the 1-$\sigma$ level. It is feasible to produce these massive galaxies with a physically realistic SFE. The stellar mass density produced by our fiducial star-formation model (with $A(z)$ chosen to match the UVLF at the corresponding redshift) in EDE is marginally consistent with findings from Wang et al. (2024a) and Desprez et al. (2024) at $z\simeq 8$ (who revisited the CEERS sample and tested removing the single most massive outlier), and the Cosmos-Web results at $z\sim 10$ (Casey et al., 2024). However, discrepancies persist at $z\simeq 12$. We explore two potential model variations, increasing $A(z)$ such that $\sigma_{\rm UV}(10^{10.5})\simeq 1.3$ mag or increasing the SFE by five times. As shown in Figure 7, these two can make the UVLF at $z\simeq 16$ consistent with observations. Since the impact of UV variability on $\rho_{\ast}$ is primarily through $\mathcal{F}(\sigma_{\rm sfr})$, it is very hard to reconcile observations purely on the UV variability direction. Alternatively, a heightened SFE emerges as a potential solution for reconciling the model with observational discrepancies in this context, as it not only increases the $M_{\ast}$ at a given $M_{\rm halo}$ but also lowers the integration range in Equation 10.
@@ Acknowledgements
We thank Robert Feldmann, and Stephanie O’Neil for suggestions that improved the manuscript. MV acknowledges support through NASA ATP 19-ATP19-0019, 19-ATP19-0020, 19-ATP19-0167, and NSF grants AST-1814053, AST-1814259, AST-1909831, AST-2007355 and AST-2107724. MBK acknowledges support from NSF CAREER award AST-1752913, NSF grants AST-1910346 and AST-2108962, NASA grant 80NSSC22K0827, and HST-GO-16686, HST-AR-17028, and HST-AR-17043 from the Space Telescope Science Institute (STScI), which is operated by AURA, Inc., under NASA contract NAS5-26555. RPN acknowledges support for this work provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by STScI.